OpenStudy (anonymous):

identify all of the following solutions of square root of x plus 14 end root plus 2 equals x the possible answers are x = -6 x = 1 x = -6 and x = 1 None of the above @StudyGurl14

OpenStudy (studygurl14):

\[\sqrt{x+14}+2=x\] \[\sqrt{x+14}=x-2\] \[x+14=(x-2)^2\] \[x +14 = x^2-4x+4\] \[0=x^2-5x-10\] Take the answer choices and plug them in and solve. If they come out to a true statement, that one is answer. If none of the answer choices work, the answer is "None of the above"

OpenStudy (studygurl14):

I'll do the first one so you can see how to do it: \[x^2-5x-10=0\] \[(-6)^2-5(-6)-10=0\] \[36 +30 -10=0\] \[56\neq0\] So -6 is not a solution. Thus, you can eliminate answer choices A and C

OpenStudy (anonymous):

is it none of the above? @StudyGurl14

OpenStudy (studygurl14):

correct

OpenStudy (anonymous):

great! I need help with a few more the next one is which of the following is a solution to 6 square root of x plus 1 equals 24 the possible answers are x = 17 x = 15 x = 5 x = 3

OpenStudy (studygurl14):

If I help you with this one will you please fan me and write a testimony?

OpenStudy (anonymous):

of course! you're helping me so much

OpenStudy (studygurl14):

Great thanks!

OpenStudy (studygurl14):

\[6\sqrt{x}+1=24\] Or... \[6\sqrt{x+1}=24\]

OpenStudy (anonymous):

is it x=3?

OpenStudy (studygurl14):

Which equation is it?

OpenStudy (anonymous):

oh sorry its the second one

OpenStudy (studygurl14):

Okay \[6\sqrt{x+1}=24\] \[\sqrt{x+1}=4\] \[x+1=4^2\] Can you do the rest yourself?

OpenStudy (anonymous):

yes thank you!

OpenStudy (studygurl14):

anytime :)

OpenStudy (anonymous):

okay two more and I promise ill leave you alone! 1. Simplify 9 fourth root of 10 end root plus 2 cubed root of 10 end root minus 6 fourth root 10 end root 6 cubed root of 10 the possible answers for this one are 3 fourth root of 10 end root 4 cubed root of 10 3 fourth root of 20 end root minus 4 cubed root of 20 negative square root of 10 negative square root of 20

OpenStudy (studygurl14):

\[9\sqrt[4]{10}+2\sqrt[3]{10}-6\sqrt[4]{10}+6\sqrt[3]{10} \] Like this?

OpenStudy (anonymous):

yes! exactly like that

OpenStudy (studygurl14):

Okay, just combine like terms. the fourth roots are like terms with each other, and the cube roots are like term each other 9 - 6 = 3 and... 2 + 6 = 8 so... \[3\sqrt[4]{10}+8\sqrt[3]{10}\]

OpenStudy (anonymous):

great thanks! and the last one is Multiply open parentheses 2 plus 5 square root of 3 close parentheses open parentheses 7 minus square root of 3 and the possible answers are 14 plus 28 square root of 3 14 minus 5 square root 3 negative 1 plus 33 square root of 3 negative 1 plus 37 square root of 3

OpenStudy (studygurl14):

\[(2+5\sqrt{3})(7-\sqrt{3})\] \[14-2\sqrt{3}+35\sqrt{3}-5(3)\] \[14+33\sqrt{3}-15\] \[-1+33\sqrt{3}\]

OpenStudy (anonymous):

thank you so so so so much!

OpenStudy (studygurl14):

You're so so so so welcome :)